OpenStudy (anonymous):

What is the circumference of a circle with a diamter of 2?

OpenStudy (asnaseer):

do you know the formula for the circumference of a circle?

OpenStudy (anonymous):

Yes, 3.14 times the radius or diameter, correct?

OpenStudy (asnaseer):

yes - which one is it - do you know?

OpenStudy (anonymous):

2pi

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Since it's the diameter I don't have to multiply by two, so just 2(3.14)?

OpenStudy (merchandize):

circumference=2 pie r we will take r=d/2 it will be2/2=1 now put the values 2*3.14*1=6.28 *pie=3.14

OpenStudy (asnaseer):

the standard formula is:\[C=2\pi r\]where C is the circumference and r is the radius. now, since the diameter is equal to twice the radius, then we can also write:\[C=2\pi r=\pi d\]where d is the diameter

OpenStudy (anonymous):

or just 6.2830 blahblahblah